# Group Decision-Making among Privacy-Aware Agents

### 15 Feb 2024 v2

@@ 2.2. Non-Private Algorithms
$\Theta^{\star}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\Lambda(\hat{\theta}),\text{ where }\Lambda(\hat{\theta})=\sum_{i=1}^{n}\log\left(\ell_{i}(\bm{S}_{i}|\hat{\theta})\right).$
@@ 2.3. Network Dynamics and Differential Privacy
The sensitivity of the joint log-likelihood is $n_{i,t}g_{i}(\hat{\theta})$ since each agent has access to $n_{i,t}$ i.i.d. observations at each point in time (in the general case) and each observation has point sensitivity $g_{i}(\hat{\theta})$.
@@ 3.1. Performance Analysis Framework
MLE (AM/GM) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{1,0,0}{\log(1/\eta)(\log(1/\eta)+\log(1/{\varepsilon}))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
MLE (Threshold) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{1,0,0}{\frac{\log(1/\eta)}{\pi^{2}}(\log(1/\eta)+\log(1/{\varepsilon})+\log(1/q))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
Online Learning $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{1,0,0}{\frac{1}{\eta}\sqrt{\frac{1}{{\varepsilon}}}}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
@@
MLE (AM/GM) $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left(n(\Gamma_{n,\Theta}+{\color[rgb]{1,0,0}{\Delta_{n,\Theta}}})\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$ $\mathcal{O}_{|\Theta|}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left({n}{\Gamma_{n,\Theta}}\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$
@@
Online Learning $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})\left(Q_{n,\Theta}+{\color[rgb]{1,0,0}{\Delta_{n,\Theta}}}\right)}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}\right)$ $\mathcal{O}_{|\Theta|}\left({\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})Q_{n,\Theta}}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}}\right)$
@@
Online Learning $Θ\mathcal{O}_{n,{\varepsilon},\eta}\left({\color[rgb]{1,0,0}{|\Theta|}}\sqrt{\frac{(Q_{n,\Theta}+{\color[rgb]{1,0,0}{|\Theta|\Delta_{n,\Theta}}})}{l_{n,\Theta}}}\right)$ $\mathcal{O}_{n}\left({\sqrt{\frac{Q_{n,\Theta}}{l_{n,\Theta}}}}\right)$
Table 2. Summary of estimators for a fixed privacy budget ${\varepsilon}>0$ and target error $\eta\in(0,1)$. We use $\mathcal{O}_{a_{1},\dots,a_{N}}\left(\cdot\right)$ to denote the big-O notation where the constants are allowed to depend on $a_{1},\dots,a_{N}$. In our algorithms, we have set the threshold $\varrho$ to be a constant. The overhead of introducing privacy is outlined in red. For compactness in notation, $\eta$ refers to $\max\{\eta_{1},\eta_{2}\}$, $\pi$ refers to $\min\{\pi_{2},\pi_{1}\}$, and $q$ refers to $\min\{1-q_{2},q_{1}\}$ (see Theorems 3.2 and 3.5). The AM/GM algorithms have matching communication complexities whenever $q=O(1)$ and $\pi=O(1/\sqrt{|\Theta|})$.
@@ 3.2. Distributed MLE
To produce the final estimates, we rely on two ways of aggregating the $K$ rounds of the algorithm: The first approach is to take the arithmetic mean of the $K$ rounds to produce the final belief, which we call the AM-Estimator. The benefit of AM-Estimator is that it can accurately w.h.p. identify all the MLE states, i.e., we can control its Type-II error probability: $\mathbb{P}\left[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}\right]\leq\eta_{2}$. This estimator has high recall but has low precision since it can recover a superset of $\Theta^{\star}$. The second way to combine the $K$ rounds is to take the geometric mean, which we call the GM-Estimator. For the GM-Estimator we can control the Type-I error probability, $\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\not\subseteq\Theta^{\star}\right]\leq\eta_{1}$, to ensure it recovers a subset of $\Theta^{\star}$ w.h.p.
@@
3.3. A Two-Threshold Algorithm for Recovering the MLE
@@
In Algorithm 4 we provide a two-threshold algorithm to simultaneously control both type-I and type-II error probabilities, with more design flexibility that comes at an increased communication cost.
@@
Algorithm 4 Private Distributed MLE (Two-Threshold Algorithm) Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta_{1},\eta_{2}$, Log-belief Threshold $\varrho>0$. Initialization: Set the number of iterations $T$ and the number of rounds $K$, and the thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ as indicated by Theorem 3.5 (for the non-asymptotic case) or Theorem 3.3 (for the asymptotic case), and $\tau=1/(1+e^{\varrho})$. Procedure: The following is repeated in $K$ rounds indexed by $k\in[K]$. In each round $k$, the agents begin by forming the noisy log-likelihoods $\bm{\sigma}_{i,k}(\hat{\theta})=e^{\bm{d}_{i,k}(\hat{\theta})}\bm{\gamma}_{i}(\hat{\theta})$, where $\bm{\gamma}_{i}(\hat{\theta})=\prod_{j=1}^{{n}_{i}}\ell_{i}(\bm{s}^{j}_{i}|\tilde{\theta})$ and $\bm{d}_{i,k}(\hat{\theta})\sim\mathcal{D}_{i}(\hat{\theta};{\varepsilon})$ independently across agents $i\in[n]$ and states $\hat{\theta}\in\Theta$. The agents initialize their beliefs to ${\bm{\nu}}_{i,k,0}(\hat{\theta})$ $=$ $\bm{\sigma}_{i,k}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,k}(\tilde{\theta})$. Over the next $T$ time steps, the agents update their belief after communicating with their neighbors, and according to the following update rule for any $\hat{\theta}$ and all $t\in[T]$: (7) ${\bm{\nu}}_{i,k,t}(\hat{\theta})=\frac{{\bm{\nu}}^{1+a_{ii}}_{i,k,t-1}(\hat{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}{\bm{\nu}}_{i,k,t-1}^{1+a_{ii}}(\tilde{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\tilde{\theta})}}\cdot$ After $K$ rounds, each agent aggregates the outcome of the rounds for every $\hat{\theta}\in\Theta$ as follows: (8) $\bm{N}_{i,T}(\hat{\theta})$ $=\frac{1}{K}\sum_{k\in[K]}\mathds{1}\left\{\bm{\nu}_{i,k,T}(\hat{\theta})>\tau\right\}\cdot$ Outputs: Return $\hat{\Theta}^{\mathrm{thres},1}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},1}\right\}\mathpunct{\raisebox{2.15277pt}{,}}\quad\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},2}\right\}\cdot$
@@
The analysis of Theorem 3.1 shows that in a given round $k$, the probability an MLE state $\theta^{\star}\in\Theta$ ends up positive as $t\to\infty$ is at least $1/|\bar{\Theta}|$, so in expectation at least $K/|\bar{\Theta}|$ rounds will yield a positive belief. On the other hand, we know that for a non-MLE state $\bar{\theta}$, on expectation, at most $(1-1/|\Theta^{\star}|)K$ trials will come up heads in expectation. For brevity, we define $p_{2}=1/|\bar{\Theta}|$ and $p_{1}=1-1/|\Theta^{\star}|$.
@@
Since $\bm{N}_{i,T}(\hat{\theta})$ is an average of independent indicator variables for all $\hat{\theta}\in\Theta$, the Chernoff bound indicates that it concentrates around its mean $\mathbb{E}\left[\bm{N}_{i,T}(\hat{\theta})\right]$. This prompts the development of the following simple algorithm, which resembles boosting algorithms such as AdaBoost (Freund and Schapire, 1997). The algorithm uses two thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ to estimate the true states as the states for which the number of times their beliefs have exceeded a threshold $\tau$ exceeds $\tau^{\mathrm{thres},2}$ and $\tau^{\mathrm{thres},1}$ respectively. The algorithm is summarized in Algorithm 4.
@@
Since $\bm{N}_{i,T}(\theta^{\star})$ for the MLE states are concentrated around a distribution with a mean that is at least $1/|\bar{\Theta}|$ and, similarly, $\bm{N}_{i,T}(\bar{\theta})$ for the non-MLE states are concentrated around their expectation which is at most $1-1/|\Theta^{\star}|$, one may rightly question whether it is possible to achieve perfect prediction of $\Theta^{\star}$ assuming the two modes are “sufficiently” well-separated. The answer to this is affirmative as long as $p_{2}>p_{1}$ and is given by the following Corollary:
@@ 3.4. Online Learning from Intermittent Streams
In Algorithm 5, we introduce the online version, where the agents have access to a data stream at every iteration. In this case, the log-belief ratio between an MLE state $\theta^{\star}$ and a non-MLE state $\bar{\theta}$ is going to converge to $(-1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)$ as $t\to\infty$ and the effect of DP noise is going to vanish as a consequence of the Césaro mean. The weak law of large numbers, and therefore the agents are going to learn the MLE states exponentially fast without the need to repeat the algorithm multiple times. The next theorem characterizes the non-asymptotic behavior of Algorithm 5:
@@
Algorithm 5 Private Online Learning Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta$, Log-belief Threshold $\varrho>0$. Initialization: Set the number of iterations $T$ as in Theorem 3.6, and $\tau=1/(1+e^{\varrho})$. Procedure: Every time $t\in\mathbb{N}_{0}$, each agent forms the likelihood product of the signals that it has received at that time period: $\bm{\gamma}_{i,t}(\hat{\theta})=\prod_{j=1}^{n_{i,t}}\ell_{i}(\bm{s}^{j}_{i,t}|\hat{\theta})$, if $n_{i,t}\geq 1$, and $\gamma_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$. The agent then draws a noise variable $\bm{d}_{i,t}(\hat{\theta})\sim\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})$ for all $i\in[n]$, $\hat{\theta}\in\Theta$ independently and forms $\bm{\sigma}_{i,t}(\hat{\theta})=e^{\bm{d}_{i,t}(\hat{\theta})}\bm{\gamma}_{i,t}(\hat{\theta})$ for all $\hat{\theta}\in\Theta$. The agent then updates its belief for $T$ iterations according to: (10) ${\bm{\nu}}_{i,t}(\hat{\theta})=\frac{\bm{\sigma}_{i,t}(\hat{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\hat{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,t}(\tilde{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\tilde{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\tilde{\theta})}}\mathpunct{\raisebox{2.15277pt}{,}}$ initialized by: ${\bm{\nu}}_{i,0}(\hat{\theta})=\bm{\sigma}_{i,0}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,0}(\tilde{\theta})$. Outputs: After $T$ iterations, return $\hat{\Theta}^{\mathrm{OL}}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{\nu}_{i,t}(\hat{\theta})\geq\tau\right\}$.

### 12 Mar 2024 v3

@@ 2.2. Non-Private Algorithms
(1) $\Theta^{\star}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\Lambda(\hat{\theta}),\text{ where }\Lambda(\hat{\theta})=\sum_{i=1}^{n}\log\left(\ell_{i}(\bm{S}_{i}|\hat{\theta})\right).$
@@ 2.3. Network Dynamics and Differential Privacy
The sensitivity of the joint log-likelihood is $n_{i,t}g_{i}(\hat{\theta})$ because each agent has access to $n_{i,t}$ i.i.d. observations at each point in time (in the general case) and each observation has point sensitivity $g_{i}(\hat{\theta})$.
@@ 3.1. Performance Analysis Framework
MLE (AM/GM) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\log(1/\eta)(\log(1/\eta)+\log(1/{\varepsilon}))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
MLE (Threshold) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\frac{\log(1/\eta)}{\pi^{2}}(\log(1/\eta)+\log(1/{\varepsilon})+\log(1/q))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
Online Learning $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\frac{1}{\eta}\sqrt{\frac{1}{{\varepsilon}}}}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
@@
MLE (AM/GM) $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left(n(\Gamma_{n,\Theta}+{\color[rgb]{0.16015625,0.5,0.7265625}{\Delta_{n,\Theta}}})\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$ $\mathcal{O}_{|\Theta|}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left({n}{\Gamma_{n,\Theta}}\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$
@@
Online Learning $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})\left(Q_{n,\Theta}+{\color[rgb]{0.16015625,0.5,0.7265625}{\Delta_{n,\Theta}}}\right)}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}\right)$ $\mathcal{O}_{|\Theta|}\left({\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})Q_{n,\Theta}}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}}\right)$
@@
Online Learning $Θ\mathcal{O}_{n,{\varepsilon},\eta}\left(\sqrt{\frac{(Q_{n,\Theta}+\Delta_{n,\Theta}){\color[rgb]{0.16015625,0.5,0.7265625}{|\Theta|}}}{l_{n,\Theta}}}\right)$ $\mathcal{O}_{n}\left({\sqrt{\frac{Q_{n,\Theta}}{l_{n,\Theta}}}}\right)$
Table 2. Summary of estimators for a fixed privacy budget ${\varepsilon}>0$ and target error $\eta\in(0,1)$. We use $\mathcal{O}_{a_{1},\dots,a_{N}}\left(\cdot\right)$ to denote the big-O notation where the constants are allowed to depend on $a_{1},\dots,a_{N}$. In our algorithms, we have set the threshold $\varrho$ to be a constant. The overhead of introducing privacy is outlined in blue. For compactness in notation, $\eta$ refers to $\max\{\eta_{1},\eta_{2}\}$, $\pi$ refers to $\min\{\pi_{2},\pi_{1}\}$, and $q$ refers to $\min\{1-q_{2},q_{1}\}$ (see Theorems 3.2 and 3.5). The AM/GM algorithms have matching communication complexities whenever $q=O(1)$ and $\pi=O(1/\sqrt{|\Theta|})$.
@@ 3.2. Distributed MLE
To produce the final estimates, we rely on two ways of aggregating the $K$ rounds of the algorithm: The first approach is to take the arithmetic mean of the $K$ rounds to produce the final belief, which we call the AM-Estimator. The benefit of AM-Estimator is that it can accurately w.h.p. identify all the MLE states, i.e., we can control its Type-II error probability: $\mathbb{P}\left[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}\right]\leq\eta_{2}$. This estimator has high recall but has low precision because it can recover a superset of $\Theta^{\star}$. The second way to combine the $K$ rounds is to take the geometric mean, which we call the GM-Estimator. For the GM-Estimator we can control the Type-I error probability, $\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\not\subseteq\Theta^{\star}\right]\leq\eta_{1}$, to ensure it recovers a subset of $\Theta^{\star}$ w.h.p.
@@
3.3. A Double-Threshold Algorithm for Recovering the MLE
@@
In Algorithm 4 we provide a double-threshold algorithm to simultaneously control both type-I and type-II error probabilities, with more design flexibility that comes at an increased communication cost.
@@
Algorithm 4 Private Distributed MLE (Double-Threshold Algorithm) Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta_{1},\eta_{2}$, Log-belief Threshold $\varrho>0$. Initialization: Set the number of iterations $T$ and the number of rounds $K$, and the thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ as indicated by Theorem 3.5 (for the non-asymptotic case) or Theorem 3.3 (for the asymptotic case), and $\tau=1/(1+e^{\varrho})$. Procedure: The following is repeated in $K$ rounds indexed by $k\in[K]$. In each round $k$, the agents begin by forming the noisy log-likelihoods $\bm{\sigma}_{i,k}(\hat{\theta})=e^{\bm{d}_{i,k}(\hat{\theta})}\bm{\gamma}_{i}(\hat{\theta})$, where $\bm{\gamma}_{i}(\hat{\theta})=\prod_{j=1}^{{n}_{i}}\ell_{i}(\bm{s}^{j}_{i}|\tilde{\theta})$ and $\bm{d}_{i,k}(\hat{\theta})\sim\mathcal{D}_{i}(\hat{\theta};{\varepsilon})$ independently across agents $i\in[n]$ and states $\hat{\theta}\in\Theta$. The agents initialize their beliefs to ${\bm{\nu}}_{i,k,0}(\hat{\theta})$ $=$ $\bm{\sigma}_{i,k}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,k}(\tilde{\theta})$. Over the next $T$ time steps, the agents update their belief after communicating with their neighbors, and according to the following update rule for any $\hat{\theta}$ and all $t\in[T]$: (8) ${\bm{\nu}}_{i,k,t}(\hat{\theta})=\frac{{\bm{\nu}}^{1+a_{ii}}_{i,k,t-1}(\hat{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}{\bm{\nu}}_{i,k,t-1}^{1+a_{ii}}(\tilde{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\tilde{\theta})}}\cdot$ After $K$ rounds, each agent aggregates the outcome of the rounds for every $\hat{\theta}\in\Theta$ as follows: (9) $\bm{N}_{i,T}(\hat{\theta})$ $=\frac{1}{K}\sum_{k\in[K]}\mathds{1}\left\{\bm{\nu}_{i,k,T}(\hat{\theta})>\tau\right\}\cdot$ Outputs: Return $\hat{\Theta}^{\mathrm{thres},1}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},1}\right\}\mathpunct{\raisebox{2.15277pt}{,}}\quad\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},2}\right\}\cdot$
@@
The analysis of Theorem 3.1 shows that in a given round $k$, the probability an MLE state $\theta^{\star}\in\Theta$ ends up positive as $t\to\infty$ is at least $1/|\bar{\Theta}|$, so in expectation at least $K/|\bar{\Theta}|$ rounds will yield a positive belief. On the other hand, we know that for a non-MLE state $\bar{\theta}$, on expectation, at most $(1-1/|\Theta^{\star}|)K$ trials will come up heads in expectation. For brevity, we define $p_{1}=1-1/|\Theta^{\star}|$ and $p_{2}=1/|\bar{\Theta}|$.
@@
Because $\bm{N}_{i,T}(\hat{\theta})$ is an average of independent indicator variables for all $\hat{\theta}\in\Theta$, the Chernoff bound indicates that it concentrates around its mean $\mathbb{E}\left[\bm{N}_{i,T}(\hat{\theta})\right]$. This prompts the development of the following simple algorithm, which resembles boosting algorithms such as AdaBoost (Freund and Schapire, 1997). The algorithm uses two thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ to estimate the true states as the states for which the number of times their beliefs have exceeded a threshold $\tau$ exceeds $\tau^{\mathrm{thres},2}$ and $\tau^{\mathrm{thres},1}$ respectively. The algorithm is summarized in Algorithm 4.
@@
Because $\bm{N}_{i,T}(\theta^{\star})$ for the MLE states are concentrated around a distribution with a mean that is at least $1/|\bar{\Theta}|$ and, similarly, $\bm{N}_{i,T}(\bar{\theta})$ for the non-MLE states are concentrated around their expectation which is at most $1-1/|\Theta^{\star}|$, one may rightly question whether it is possible to achieve perfect prediction of $\Theta^{\star}$ assuming the two modes are “sufficiently” well-separated. The answer to this is affirmative as long as $p_{2}>p_{1}$ and is given by the following Corollary:
@@
  • The optimal distributions that minimize the convergence time $T$ for both estimators, are $\mathcal{D}_{i}^{\star}({\varepsilon})=\mathrm{Lap}\left({\Delta_{i}K|\Theta|}/{{\varepsilon}}\right)$.
@@ 3.4. Online Learning from Intermittent Streams
In Algorithm 5, we introduce the online learning version, where the agents have access to a data stream at every iteration. In this algorithm, the agents receive $n_{i,t}\geq 1$ signals at every iteration $t$, calculate the likelihood and the noisy version of it (in order to preserve DP) and then aggregate it with their self- and neighboring beliefs from iteration $t-1$. Here, the data are produced by a model with true parameters $\theta^{\circ}\in\Theta$, and the set of identified states is given as the singleton
$\theta^{\star}=\theta^{\circ}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\;\mathbb{E}_{\theta^{\circ}}\left[\Lambda(\hat{\theta})\right],$
for any true model $\theta^{\circ}\in\Theta$, assuming a collectively identifiable statistical model; collective in the sense that $\Lambda(\hat{\theta})$ in Equation 1 is defined as the sum of the individual log-likelihoods.
@@
Algorithm 5 Private Online Learning Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta$. Initialization: Set the number of iterations $T$ as in Theorem 3.7. Procedure: Every time $t\in\mathbb{N}_{0}$, each agent forms the likelihood product of the signals that it has received at that time period: $\bm{\gamma}_{i,t}(\hat{\theta})=\prod_{j=1}^{n_{i,t}}\ell_{i}(\bm{s}^{j}_{i,t}|\hat{\theta})$, if $n_{i,t}\geq 1$, and $\gamma_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$. The agent then draws a noise variable $\bm{d}_{i,t}(\hat{\theta})\sim\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})$ for all $i\in[n]$, $\hat{\theta}\in\Theta$ independently and forms $\bm{\sigma}_{i,t}(\hat{\theta})=e^{\bm{d}_{i,t}(\hat{\theta})}\bm{\gamma}_{i,t}(\hat{\theta})$ for all $\hat{\theta}\in\Theta$ (and set $\bm{\sigma}_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$). The agent then updates its belief for $T$ iterations according to: (11) ${\bm{\nu}}_{i,t}(\hat{\theta})=\frac{\bm{\sigma}_{i,t}(\hat{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\hat{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,t}(\tilde{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\tilde{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\tilde{\theta})}}\mathpunct{\raisebox{2.15277pt}{,}}$ initialized by: ${\bm{\nu}}_{i,0}(\hat{\theta})=\bm{\sigma}_{i,0}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,0}(\tilde{\theta})$. Outputs: After $T$ iterations, return $\hat{\theta}^{\mathrm{OL}}_{i,T}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\bm{\nu}_{i,t}(\hat{\theta})$.
For the asymptotic regime, the log-belief ratio between $\theta^{\star}$ and a non-MLE state $\bar{\theta}$ is going to converge to $(-1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)$ as $t\to\infty$ and the effect of DP noise is going to vanish as a consequence of the Césaro mean. The weak law of large numbers, and therefore the agents are going to learn the MLE states exponentially fast without the need to repeat the algorithm multiple times. For completeness, we state the asymptotic theorem and provide a proof sketch:
Theorem 3.6 ().
As $T\to\infty$, for any distributions $\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})=\mathcal{D}_{i,t}({\varepsilon})$ that satisfy ${\varepsilon}$-DP and do not depend on the state $\hat{\theta}$, Algorithm 5, we have that: If $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$ for all $\bar{\theta}\neq\theta^{\star}$, then $\lim_{T\to\infty}\bm{\nu}_{i,t}(\theta^{\star})=1$ and the agents learn $\theta^{\star}$ asymptotically fast with rate $-l_{n,\Theta}/n$.
Proof Sketch. We can show that the log-belief ratio $\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})$ between a pair of states $\hat{\theta}\in\Theta$ and $\check{\theta}\in\Theta$ obeys $\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})=\sum_{\tau=0}^{t}A^{\tau}\left(\overline{\bm{\lambda}}_{t-\tau}(\hat{\theta},\check{\theta})+\overline{\bm{d}}_{t-\tau}(\hat{\theta})-\overline{\bm{d}}_{t-\tau}(\check{\theta})\right)$ where $\overline{\bm{\lambda}}_{t-\tau}(\hat{\theta},\check{\theta})$ is the log-likelihood ratio between $\hat{\theta}$ and $\check{\theta}$ at iteration $t-\tau$ and $\overline{\bm{d}}_{t-\tau}(\hat{\theta}),\overline{\bm{d}}_{t-\tau}(\check{\theta})$ are the vectorized noise variables respectively. The Césaro mean together with the strong law of large numbers implies that $\lim_{t\to\infty}(1/t)\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})=(1/n)\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})+\overline{\bm{d}}_{0}(\hat{\theta})-\overline{\bm{d}}_{0}(\check{\theta})\right]\mathds{1}=(1/n)\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})\right]\mathds{1}$, as $\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})-\overline{\bm{d}}_{0}(\check{\theta})\right]\mathds{1}=\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})\right]\mathds{1}-\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})\right]\mathds{1}=0$ because the noise does not depend on the states. Moreover, $\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})\right]\mathds{1}=\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)\mathds{1}$, and whenever $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$, the agents can learn the true state exponentially fast with rate $-l_{n,\Theta}/n$. ∎
The next theorem characterizes the non-asymptotic behavior of Algorithm 5: